OpenStudy (anonymous):

Exact value of : cot(15 degrees) sin(195 degrees) cos(7pi/12) Only fractions and radicals. No decimals.

OpenStudy (anonymous):

@divu.mkr

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (kc_kennylau):

Hint: 15=45-30, 195=180+15, 7pi/12=pi/2+pi/12

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (kc_kennylau):

Yeah, go on calling people to help you

OpenStudy (kc_kennylau):

Keep being lazy by not using your brain

OpenStudy (anonymous):

ha you don't even know half of the work I do

OpenStudy (anonymous):

I've got a 4.5 gpa just that trig isn't my strong point and you know what I'm kinda tired its the end of the week

OpenStudy (kc_kennylau):

Formulas that will be useful: \[\sin(A\pm B)=\sin A\cos B\pm\cos A\sin B\]\[\cos(A\pm B)=\cos A\cos B\mp\sin A\sin B\]\[\tan(A\pm B)=\frac{\tan A\pm\tan B}{1\mp\tan A\tan B}\]

OpenStudy (anonymous):

thank you I already have them but don't understand how to apply them

OpenStudy (anonymous):

what would be A? B?

OpenStudy (kc_kennylau):

A and B can be anything you like, and the equality still stands.

OpenStudy (kc_kennylau):

For example you can put A=45deg and B=30deg to the first one, you'll get: \[\sin(45^\circ+30^\circ)=\sin45^\circ\cos30^\circ+\cos45^\circ\sin30^\circ\]

OpenStudy (kc_kennylau):

Simplify that and you'll get \(\displaystyle\sin(75^\circ)=\frac1{\sqrt2}\cdot\frac{\sqrt3}2+\frac1{\sqrt2}\cdot\frac12\)

OpenStudy (kc_kennylau):

Simplify that further and you'll have the value of \(\sin75^\circ\), get it?

OpenStudy (anonymous):

great

OpenStudy (kc_kennylau):

Yes or no

OpenStudy (kc_kennylau):

You have to give me response or I can't help you

OpenStudy (anonymous):

yeah

OpenStudy (kc_kennylau):

Good, now find the value of \(\tan15^\circ\) by using the formula

OpenStudy (anonymous):

okayy well I already got the answers to all of them could you check them?

OpenStudy (kc_kennylau):

Okay

OpenStudy (kc_kennylau):

What are your answers?

OpenStudy (anonymous):

cot 15 = 2 + radical 3 sin 195 = - (radical 3 - 1) / 2 radical2 cos 7pi/12 = -(radical 3 - 1) / 2 radical2

OpenStudy (kc_kennylau):

All correct

OpenStudy (anonymous):

Awesome thanks!